# Pre-Generating Multi-Difficulty PDE Data for Few-Shot Neural PDE Solvers

### v1

@@ C.3 Poseidon-T
We evaluate Poseidon-T using the pretrained model checkpoints provided by herde2024poseidon, available on Hugging Face. [^3] https://huggingface.co/camlab-ethz/Poseidon-T We perform fine-tuning on our custom datasets without any further pretraining.
@@ C.4 Poseidon-B
We evaluate Poseidon-B using the pretrained model checkpoint provided by herde2024poseidon, available on Hugging Face. [^4] https://huggingface.co/camlab-ethz/Poseidon-B We perform fine-tuning on our custom datasets without any further pretraining. The architecture mirrors Poseidon-T’s setup, following a SwinUNet-based transformer backbone with hierarchical (U-Net style) multiscale attention and window-based self-attention. We retain the pretrained model structure and update weights via supervised autoregressive fine-tuning. Key configuration details include:
@@ C.5 Poseidon-L
We evaluate Poseidon-L using the pretrained model checkpoint provided by herde2024poseidon, available on Hugging Face. [^6] https://huggingface.co/camlab-ethz/Poseidon-L We fine-tune this model on our custom datasets with no additional pretraining. The architecture is identical to the other Poseidon variants, employing the same SwinUNet-style transformer backbone with hierarchical multiscale attention and window-based (shifted-window) self-attention. We preserve the original model architecture and learn weights via supervised autoregressive fine-tuning. Key configuration details include:

### v2

@@ C.3 Poseidon-T
We evaluate Poseidon-T using the pretrained model checkpoints provided by Herde et al. [2024], available on Hugging Face. [^3] https://huggingface.co/camlab-ethz/Poseidon-T We perform fine-tuning on our custom datasets without any further pretraining.
@@ C.4 Poseidon-B
We evaluate Poseidon-B using the pretrained model checkpoint provided by Herde et al. [2024], available on Hugging Face. [^4] https://huggingface.co/camlab-ethz/Poseidon-B We perform fine-tuning on our custom datasets without any further pretraining. The architecture mirrors Poseidon-T’s setup, following a SwinUNet-based transformer backbone with hierarchical (U-Net style) multiscale attention and window-based self-attention. We retain the pretrained model structure and update weights via supervised autoregressive fine-tuning. Key configuration details include:
@@ C.5 Poseidon-L
We evaluate Poseidon-L using the pretrained model checkpoint provided by Herde et al. [2024], available on Hugging Face. [^6] https://huggingface.co/camlab-ethz/Poseidon-L We fine-tune this model on our custom datasets with no additional pretraining. The architecture is identical to the other Poseidon variants, employing the same SwinUNet-style transformer backbone with hierarchical multiscale attention and window-based (shifted-window) self-attention. We preserve the original model architecture and learn weights via supervised autoregressive fine-tuning. Key configuration details include: